# Gravity as a topological gauge theory

### v3

@@ 2.1 Lorentzian (dS and AdS) spacetimes as mutations of an
where $J_{ab}$ is canonically identified with the matrix representation of $\mathfrak{h}\subset\mathfrak{g}$, $\varphi^{a}_{\mathfrak{m}}=\varphi^{a}_{\mathfrak{p}}$ and $M_{a}=\begin{pmatrix}0&e_{a}\\ k\bar{e}_{a}&0\end{pmatrix}$ the generators of $\mathfrak{m}$.
Let $\varphi,\varphi^{\prime}\in\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}$, their Lie bracket is given by the commutator and reads:
$[\varphi,\varphi^{\prime}]$ $=\begin{bmatrix}\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ k\varphi_{\bar{\mathfrak{p}}}&0\end{pmatrix},\begin{pmatrix}\varphi^{\prime}_{\mathfrak{h}}&\varphi^{\prime}_{\mathfrak{p}}\\ k\varphi^{\prime}_{\bar{\mathfrak{p}}}&0\end{pmatrix}\end{bmatrix}=\begin{pmatrix}[\varphi_{\mathfrak{h}},\varphi^{\prime}_{h}]+k(\varphi_{\mathfrak{p}}\varphi^{\prime}_{\bar{\mathfrak{p}}}-\varphi^{\prime}_{\mathfrak{p}}\varphi_{\bar{\mathfrak{p}}})&(\varphi_{\mathfrak{h}}\varphi^{\prime}_{\mathfrak{p}}-\varphi^{\prime}_{\mathfrak{h}}\varphi_{\mathfrak{p}})\\ k(\varphi_{\bar{\mathfrak{p}}}\varphi^{\prime}_{\mathfrak{h}}-\varphi^{\prime}_{\bar{\mathfrak{p}}}\varphi_{\mathfrak{h}})&0\end{pmatrix}$
$=\bigl{(}\dfrac{1}{8}\varphi_{\mathfrak{h}}^{ab}\varphi_{h}^{\prime cd}C_{ab,cd}^{ef}+\dfrac{k}{2}(\varphi_{\mathfrak{p}}^{e}\varphi_{\mathfrak{p}}^{\prime f}-\varphi_{\mathfrak{p}}^{f}\varphi_{\mathfrak{p}}^{\prime e})\bigr{)}J_{ef}\oplus\dfrac{1}{2}(\varphi_{\mathfrak{h}}^{ab}\varphi_{\mathfrak{p}}^{\prime c}-\varphi_{\mathfrak{h}}^{\prime ab}\varphi_{\mathfrak{p}}^{c})(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})M_{r}$ (8)
where we have used the commutator $[J_{ab},M_{c}]=\eta_{bc}M_{a}-\eta_{ac}M_{b}$.
A Cartan connection $\varpi$ with this $G/H$ Cartan geometry and this mutation can be decomposed as:
$\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{m}}=\begin{pmatrix}\dfrac{1}{2}\varpi_{\mathfrak{h}}^{ab}J_{ab}&\varpi_{\mathfrak{p}}^{a}e_{a}\\ k\varpi_{\mathfrak{p}}^{a}\bar{e}_{a}&0\end{pmatrix}=\begin{pmatrix}\dfrac{1}{2}A^{ab}J_{ab}&\dfrac{1}{\ell}\beta^{a}e_{a}\\ \dfrac{k}{\ell}\beta^{a}\bar{e}_{a}&0\end{pmatrix}=\begin{pmatrix}A&\dfrac{1}{\ell}\beta\\ \dfrac{k}{\ell}\bar{\beta}&0\end{pmatrix}$ (9)
where $\varpi_{\mathfrak{h}}^{ab}=\varpi^{ab}_{\mathfrak{h}\>\,\mu}dx^{\mu}=A^{ab}_{~{}\,\,\mu}dx^{\mu}$ corresponds to the spin connection and $\ell\varpi_{\mathfrak{p}}^{a}=\beta^{\,a}_{~{}\,\,\mu}dx^{\mu}$ corresponds to the tetrad while $\varpi_{\mathfrak{p}}=\dfrac{1}{\ell}\beta$ is the soldering form and $\bar{\beta}=\beta^{\text{T}}\eta$.
@@ A Cartan connection $\varpi$ with this $G/H$ Cartan geometry and this mutation can be decomposed as:
$\bar{\Omega}$ $=d\varpi+\dfrac{1}{2}[\varpi,\varpi]=\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{m}}$ (10)
$=\bigl{(}d\varpi_{\mathfrak{h}}+\dfrac{1}{2}[\varpi_{\mathfrak{h}},\varpi_{\mathfrak{h}}]+k\xi\bigr{)}$ (11)
$\qquad\oplus\bigl{(}d\varpi_{\mathfrak{p}}+\dfrac{1}{4}(\varpi^{ab}_{\mathfrak{h}\>\,\mu}\varpi^{~{}\,c}_{\mathfrak{p}\,\,\nu}-\varpi^{ab}_{\mathfrak{h}\>\,\nu}\varpi^{~{}\,c}_{\mathfrak{p}\,\,\mu})(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})dx^{\mu}\wedge dx^{\nu}\bigr{)}\otimes M_{r}$
@@
$\qquad\oplus\dfrac{1}{\ell}\bigl{(}\partial_{\mu}\beta^{\,r}_{~{}\,\,\nu}+\dfrac{1}{2}A^{ab}_{~{}\,\,\mu}\beta^{\,c}_{~{}\,\,\nu}(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})\bigr{)}dx^{\mu}\wedge dx^{\nu}\otimes M_{r}$ (12)
We identify the curvature and torsion of Einstein-Cartan gravity:
$R$ $=\bar{\Omega}_{\mathfrak{h}}-k\xi=dA+\dfrac{1}{2}[A,A]$ (13)
$T$ $=\ell\bar{\Omega}_{\mathfrak{m}}=d\beta+A\wedge\beta=\bigl{(}\partial_{\mu}\beta^{\,r}_{~{}\,\,\nu}+\dfrac{1}{2}A^{ab}_{~{}\,\,\mu}\beta^{\,c}_{~{}\,\,\nu}(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})\bigr{)}dx^{\mu}\wedge dx^{\nu}\otimes M_{r}$ (14)
with $\xi=\dfrac{1}{2}\xi^{ab}J_{ab}=\dfrac{1}{4}\xi^{ab}_{~{}~{}\mu\nu}dx^{\mu}\wedge dx^{\nu}\otimes J_{ab}$, where $\xi^{ab}=2\varpi_{\mathfrak{p}}^{a}\wedge\varpi_{\mathfrak{p}}^{b}$ and $\xi^{ab}_{~{}~{}\mu\nu}=2(\varpi^{~{}\,a}_{\mathfrak{p}\,\,\mu}\varpi^{~{}\,b}_{\mathfrak{p}\,\,\nu}-\varpi^{~{}\,a}_{\mathfrak{p}\,\,\nu}\varpi^{~{}\,b}_{\mathfrak{p}\,\,\mu})$ an additional term due to the symmetric algebra structure.
The Bianchi identities $D\bar{\Omega}=d\bar{\Omega}+[\varpi,\bar{\Omega}]=0$ split as:
@@
with $\bar{T}=T^{\text{T}}\eta$.
2.2 Hodge operator
The Hodge operator defined on differential forms on the manifold $\mathscr{M}$ requires a metric $g$ on $\mathscr{M}$. According to the reductive Cartan geometry $G/H$ at hand, $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, [^6] In this Section, one denotes $\mathfrak{g}$ the mutated algebra $\mathfrak{g}^{\prime}$ defined in (1) and $\mathfrak{p}\simeq\mathfrak{m}$. the Killing form $K$ on $\mathfrak{g}$ splits into two parts:
@@ 2.2 Hodge operator
Let $\kappa$ and $\zeta$ be real scalars. In this case we choose $h=\kappa K$ as the metric on the Lie algebra $\mathfrak{g}$ and $g=\zeta\varpi_{\mathfrak{m}}^{*}h=\kappa\zeta\varpi_{\mathfrak{m}}^{*}K_{\mathfrak{m}}$. It corresponds to the pullback of the $\mathfrak{m}$ part of the Killing form on $\mathfrak{g}$ by $\varpi_{\mathfrak{m}}$ the $\mathfrak{m}$-part of the Cartan connection up to the factor $\kappa\zeta$.
For a manifold $\mathscr{M}$ of dimension $m=4$ with Lorentzian Cartan geometry defined by the mutation (1):
$K_{\mathfrak{g}}(\varphi,\varphi^{\prime})$ $=K_{\mathfrak{h}}(\varphi_{\mathfrak{h}},\varphi^{\prime}_{\mathfrak{h}})+K_{\mathfrak{m}}(\varphi_{\mathfrak{m}},\varphi^{\prime}_{\mathfrak{m}})=3\operatorname{Tr}(\varphi_{\mathfrak{h}},\varphi^{\prime}_{\mathfrak{h}})+6k\eta(\varphi_{\mathfrak{m}},\varphi^{\prime}_{\mathfrak{m}})$
$=-3\eta_{ac}\eta_{bd}\varphi_{\mathfrak{h}}^{ab}\varphi_{\mathfrak{h}}^{\prime cd}+6k\eta_{ab}\varphi_{\mathfrak{m}}^{a}\varphi_{\mathfrak{m}}^{\prime b}$
with $K_{\mathfrak{m},ab}=6k\eta_{ab}$ and thus
@@
Let $\omega\in\Omega^{r}(\mathscr{M},\mathfrak{g})$ be an $r$-form on $\mathscr{M}$ with values in $\mathfrak{g}$. Its local trivialisation in a given chart is :
@@
Let $\varepsilon_{a_{1}a_{2}...a_{n}}$ be the Levi-Civita symbol in dimension $n$, such that $\varepsilon_{12...n}=1$. From the metric $g$ one can define a Hodge star operator * that acts on $\omega$ as :
$*\omega=$ $\dfrac{1}{r!}\sqrt{|det(g)|}\,\omega_{\mu_{1}...\mu_{r}}g^{\mu_{1}\nu_{1}}...g^{\mu_{r}\nu_{r}}\varepsilon_{\nu_{1}...\nu_{m}}dx^{\nu_{r+1}}\wedge...dx^{\nu_{m}}$ (3)

### v5

@@ 2.1 Lorentzian (dS and AdS) spacetimes as mutations of an
where $J_{ab}$ is canonically identified with the matrix representation $J_{ab}=\begin{pmatrix}J_{ab}&0\\ 0&0\end{pmatrix}$ of $\mathfrak{h}\subset\mathfrak{g}^{\prime}$, $\varphi^{a}_{\mathfrak{m}}=\varphi^{a}_{\mathfrak{p}}$ and $M_{a}=\begin{pmatrix}0&k_{1}e_{a}\\ k_{2}\bar{e}_{a}&0\end{pmatrix}=\mu\big{(}\begin{pmatrix}0&e_{a}\\ 0&0\end{pmatrix}\big{)}$ are the generators of $\mathfrak{m}$ constructed out of the canonical basis of $\mathfrak{p}=\mathfrak{\mathbb{R}}^{3,1}$ and the parameters of the mutation.
Let $\varphi,\varphi^{\prime}\in\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}$, their Lie bracket is given by the commutator:
$[\varphi,\varphi^{\prime}]$ $=\begin{bmatrix}\begin{pmatrix}\varphi_{\mathfrak{h}}&k_{1}\varphi_{\mathfrak{p}}\\ k_{2}\bar{\varphi}_{\mathfrak{p}}&0\end{pmatrix},\begin{pmatrix}\varphi^{\prime}_{\mathfrak{h}}&k_{1}\varphi^{\prime}_{\mathfrak{p}}\\ k_{2}\bar{\varphi}^{\prime}_{\mathfrak{p}}&0\end{pmatrix}\end{bmatrix}=\begin{pmatrix}[\varphi_{\mathfrak{h}},\varphi^{\prime}_{h}]+k_{1}k_{2}(\varphi_{\mathfrak{p}}\bar{\varphi}^{\prime}_{\mathfrak{p}}-\varphi^{\prime}_{\mathfrak{p}}\bar{\varphi}_{\mathfrak{p}})&k_{1}(\varphi_{\mathfrak{h}}\varphi^{\prime}_{\mathfrak{p}}-\varphi^{\prime}_{\mathfrak{h}}\varphi_{\mathfrak{p}})\\ k_{2}(\bar{\varphi}_{\mathfrak{p}}\varphi^{\prime}_{\mathfrak{h}}-\bar{\varphi}^{\prime}_{\mathfrak{p}}\varphi_{\mathfrak{h}})&0\end{pmatrix}$
$=\bigl{(}\dfrac{1}{8}\varphi_{\mathfrak{h}}^{ab}\varphi_{h}^{\prime cd}C_{ab,cd}^{ef}+\dfrac{k_{1}k_{2}}{2}(\varphi_{\mathfrak{p}}^{e}\varphi_{\mathfrak{p}}^{\prime f}-\varphi_{\mathfrak{p}}^{f}\varphi_{\mathfrak{p}}^{\prime e})\bigr{)}J_{ef}\oplus\dfrac{1}{2}(\varphi_{\mathfrak{h}}^{ab}\varphi_{\mathfrak{p}}^{\prime c}-\varphi_{\mathfrak{h}}^{\prime ab}\varphi_{\mathfrak{p}}^{c})(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})M_{r}$
where we have used the commutators $[J_{ab},M_{c}]=\eta_{bc}M_{a}-\eta_{ac}M_{b}$ and $[M_{a},M_{b}]=k_{1}k_{2}\,J_{ab}$. [^7] It is interesting to notice that the product of the mutation parameters $k_{1}$ and $k_{2}$ can be related to a deformation parameter $\alpha$ as given in [15, eq.(2.1)].
In order to link our notation with the one used in [1], we identify $k_{1}=\dfrac{1}{\ell}$ and $k_{2}=\dfrac{k}{\ell}$. A Cartan connection $\varpi$ associated to this mutated $G/H$ Cartan geometry can be decomposed as:
$\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{m}}=\begin{pmatrix}\dfrac{1}{2}\varpi_{\mathfrak{h}}^{ab}J_{ab}&\dfrac{1}{\ell}\varpi_{\mathfrak{p}}^{a}e_{a}\\ \dfrac{k}{\ell}\varpi_{\mathfrak{p}}^{a}\bar{e}_{a}&0\end{pmatrix}=\begin{pmatrix}\dfrac{1}{2}A^{ab}J_{ab}&\dfrac{1}{\ell}\beta^{a}e_{a}\\ \dfrac{k}{\ell}\beta^{a}\bar{e}_{a}&0\end{pmatrix}=\begin{pmatrix}A&\dfrac{1}{\ell}\beta\\ \dfrac{k}{\ell}\bar{\beta}&0\end{pmatrix}$ (8)
where $\varpi_{\mathfrak{h}}^{ab}=\varpi^{ab}_{\mathfrak{h}\>\,\mu}dx^{\mu}=A^{ab}_{~{}\,\,\mu}dx^{\mu}$ corresponds to the spin connection and $\varpi_{\mathfrak{m}}=\beta^{a}M_{a}$ is the $\mathfrak{g}^{\prime}/\mathfrak{h}=\mathfrak{m}$-valued soldering form (see [11, Chap.5, §3, Definition 3.1]) after mutation.
@@
$\bar{\Omega}$ $=d\varpi+\dfrac{1}{2}[\varpi,\varpi]=\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{m}}=\begin{pmatrix}R+\dfrac{k}{\ell^{2}}\beta\wedge\bar{\beta}&0\\ 0&0\end{pmatrix}+\dfrac{1}{\ell}\begin{pmatrix}0&T\\ k\bar{T}&0\end{pmatrix}$ (9)
$=\bigl{(}d\varpi_{\mathfrak{h}}+\dfrac{1}{2}[\varpi_{\mathfrak{h}},\varpi_{\mathfrak{h}}]+\dfrac{1}{2}[\varpi_{\mathfrak{m}},\varpi_{\mathfrak{m}}]\bigr{)}$ (10)
$\qquad\oplus\bigl{(}d\varpi_{\mathfrak{p}}^{r}+\dfrac{1}{4}(\varpi^{ab}_{\mathfrak{h}\>\,\mu}\varpi^{~{}\,c}_{\mathfrak{p}\,\,\nu}-\varpi^{ab}_{\mathfrak{h}\>\,\nu}\varpi^{~{}\,c}_{\mathfrak{p}\,\,\mu})(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})dx^{\mu}\wedge dx^{\nu}\bigr{)}\otimes M_{r}$
@@
$\qquad\oplus\bigl{(}\partial_{\mu}\beta^{\,r}_{~{}\,\,\nu}+\dfrac{1}{2}A^{ab}_{~{}\,\,\mu}\beta^{\,c}_{~{}\,\,\nu}(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})\bigr{)}dx^{\mu}\wedge dx^{\nu}\otimes M_{r}$ (11)
with $\bar{\Omega}_{\mathfrak{m}}$ the torsion of the Cartan connection $\varpi$ with values in $\mathfrak{g}^{\prime}/\mathfrak{h}=\mathfrak{m}$. For further use it is worthwhile to notice that the graded bracket
$\tfrac{1}{2}[\varpi_{\mathfrak{m}},\varpi_{\mathfrak{m}}]=\dfrac{k}{\ell^{2}}\begin{pmatrix}\beta\wedge\bar{\beta}&0\\ 0&0\end{pmatrix}=\begin{pmatrix}k\,\xi&0\\ 0&0\end{pmatrix}\in\mathfrak{h}$ (12)
with $\xi=\dfrac{1}{\ell^{2}}\beta\wedge\bar{\beta}=\dfrac{1}{2}\xi^{ab}J_{ab}=\dfrac{1}{4}\xi^{ab}_{~{}~{}\mu\nu}dx^{\mu}\wedge dx^{\nu}\otimes J_{ab}$, where $\xi^{ab}=\dfrac{1}{\ell^{2}}\varpi_{\mathfrak{p}}^{a}\wedge\varpi_{\mathfrak{p}}^{b}$ and $\xi^{ab}_{~{}~{}\mu\nu}=\dfrac{1}{\ell^{2}}(\varpi^{~{}\,a}_{\mathfrak{p}\,\,\mu}\varpi^{~{}\,b}_{\mathfrak{p}\,\,\nu}-\varpi^{~{}\,a}_{\mathfrak{p}\,\,\nu}\varpi^{~{}\,b}_{\mathfrak{p}\,\,\mu})$ an additional term due to the symmetric Lie algebra structure. Let us add that (9) and (12) exemplify [11, Proposition 6.3, p.218].
We also identify the spacetime curvature and torsion of Einstein-Cartan gravity:
$R$ $=\bar{\Omega}_{\mathfrak{h}}-k\xi=dA+\dfrac{1}{2}[A,A]=dA+A^{2}$ (13)
$T$ $=d\beta+A\wedge\beta=T^{a}e_{a}=\bigl{(}\partial_{\mu}\beta^{\,r}_{~{}\,\,\nu}+\dfrac{1}{2}A^{ab}_{~{}\,\,\mu}\beta^{\,c}_{~{}\,\,\nu}(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})\bigr{)}dx^{\mu}\wedge dx^{\nu}\otimes e_{r}$ (14)
The Bianchi identity $D\bar{\Omega}=d\bar{\Omega}+[\varpi,\bar{\Omega}]=0$ splits as:
@@
with $\bar{T}=d\bar{\beta}+\bar{\beta}\wedge A=T^{a}\bar{e}_{a}=\eta(T)$.
2.2 Hodge *-operator
The Hodge operator defined on differential forms on the manifold $\mathscr{M}$ requires a metric $g$ on $\mathscr{M}$. According to the reductive Cartan geometry $G/H$ at hand, $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, [^8] In this Section, one denotes $\mathfrak{g}$ the mutated algebra $\mathfrak{g}^{\prime}$ defined in (1) and $\mathfrak{p}\simeq\mathfrak{m}$. the Killing form $K_{\mathfrak{g}}(\varphi,\varphi^{\prime})=\operatorname{Tr}(ad_{\varphi}\circ ad_{\varphi^{\prime}})$ on $\mathfrak{g}$ splits into two parts:
@@ 2.2 Hodge *-operator
where the trace is performed with respect to the generators $J_{ab},a<b$ and $M_{a}$. [^9] In full generality, one has $\operatorname{Tr}(ad_{\varphi}\circ ad_{\varphi^{\prime}})=(m-1)\operatorname{Tr}(\varphi_{\mathfrak{h}}\varphi^{\prime}_{\mathfrak{h}})+2k_{1}k_{2}(m-1)\eta(\varphi_{\mathfrak{p}},\varphi^{\prime}_{\mathfrak{p}})$.
Let $\kappa$ and $\zeta$ be real scalars. In this case we choose $h=\kappa K_{\mathfrak{g}}$ as the metric on the Lie algebra $\mathfrak{g}$ and $g=\zeta\varpi_{\mathfrak{m}}^{*}h=\kappa\zeta\varpi_{\mathfrak{m}}^{*}K_{\mathfrak{m}}$. It corresponds to the pullback of the $\mathfrak{m}$-part of the Killing form on $\mathfrak{g}$ by $\varpi_{\mathfrak{m}}$ the $\mathfrak{m}$-part of the Cartan connection up to the factor $\kappa\zeta$.
For a manifold $\mathscr{M}$ of dimension $m=4$ with Lorentzian Cartan geometry defined by the mutation (1) with $k_{1}=1/\ell$ and $k_{2}=k/\ell$:
$K_{\mathfrak{g}}(\varphi,\varphi^{\prime})$ $=K_{\mathfrak{h}}(\varphi_{\mathfrak{h}},\varphi^{\prime}_{\mathfrak{h}})+K_{\mathfrak{m}}(\varphi_{\mathfrak{m}},\varphi^{\prime}_{\mathfrak{m}})=3\operatorname{Tr}(\varphi_{\mathfrak{h}}\varphi^{\prime}_{\mathfrak{h}})+\dfrac{6k}{\ell^{2}}\eta(\varphi_{\mathfrak{p}},\varphi^{\prime}_{\mathfrak{p}})$
$=-3\eta_{ac}\eta_{bd}\varphi_{\mathfrak{h}}^{ab}\varphi_{\mathfrak{h}}^{\prime cd}+\dfrac{6k}{\ell^{2}}\eta_{ab}\varphi_{\mathfrak{m}}^{a}\varphi_{\mathfrak{m}}^{\prime b}$ (2)
with $K_{\mathfrak{m}}(M_{a},M_{b})=K_{\mathfrak{m},ab}=\dfrac{6k}{\ell^{2}}\eta_{ab}$ and thus, for any vector fields $X,Y$ on $\mathscr{M}$,
@@
Let $\omega\in\Omega^{r}(\mathscr{M},\mathfrak{g})$ be an $r$-form on $\mathscr{M}$ with values in $\mathfrak{g}$. Its local trivialization in a given chart is:
@@
Let $\varepsilon_{a_{1}a_{2}...a_{m}}$ be the Levi-Civita symbol in dimension $m$, such that $\varepsilon_{12...m}=1$. From the metric $g$ and its determinant $|g|$, one can define a Hodge star operator * that acts on $\omega$ as:
$*\omega=$ $\dfrac{1}{r!}\sqrt{|g|}\,\omega_{\mu_{1}...\mu_{r}}g^{\mu_{1}\nu_{1}}...g^{\mu_{r}\nu_{r}}\varepsilon_{\nu_{1}...\nu_{m}}dx^{\nu_{r+1}}\wedge...dx^{\nu_{m}}$ (4)